# LeTO: Learning Constrained Visuomotor Policy with Differentiable Trajectory Optimization

### v2

@@ Abstract
This paper introduces LeTO, a method for learning constrained visuomotor policy via differentiable trajectory optimization. Our approach integrates a differentiable optimization layer into the neural network. By formulating the optimization layer as a trajectory optimization problem, we enable the model to end-to-end generate actions in a safe and constraint-controlled fashion without extra modules. Our method allows for the introduction of constraints information during the training process, thereby balancing the training objectives of satisfying constraints, smoothing the trajectories, and minimizing errors with demonstrations. This “gray box” method marries the optimization-based safety and interpretability with the powerful representational abilities of neural networks. We quantitatively evaluate LeTO in simulation and on the real robot. In simulation, LeTO achieves a success rate comparable to state-of-the-art imitation learning methods, but the generated trajectories are of less uncertainty, higher quality, and smoother. In real-world experiments, we deployed LeTO to handle constraints-critical tasks. The results show the effectiveness of LeTO comparing with state-of-the-art imitation learning approaches. We release our code at https://github.com/ZhengtongXu/LeTO.
@@ I Introduction
Imitation learning [1] focuses on deriving robot policies from demonstrations. This process can be formulated as a supervised learning task, aiming to learn the mapping between observations and robot actions. Through imitation learning, robots can perform highly complex and diverse tasks [2, 3, 4, 5, 6, 7, 8].
However, ensuring the safety of robots is very important due to the interaction between robots and the real world. In contrast to optimization-based trajectory generation methods, imitation learning often exhibits greater uncertainty, which can lead to system instability, reduced robustness, and safety concerns. Hence, in comparison to other supervised learning problems, this pose novel and exceptional challenges in safety for imitation learning.
Existing works focus on different aspects and challenges of imitation learning, such as improving action accuracy by addressing compounding errors [4] and represent multi-modal distributions by implicit policy [9, 10, 2]. However, these methods overlook the critical aspects of safety and smoothness of generated trajectories. For robotic systems, their behaviors need to meet specific constraints to ensure the safety of the system. Moreover, for certain tasks, the robot actions also need to meet specific constraints to ensure their successful completion. In addition, adding explainable constraints for neural network-based algorithms, often called “black boxes”, is much harder than for traditional model-based methods.
@@
1) Differentiable optimization layer: We integrate a differentiable optimization layer to the neural network and formulate it as a trajectory optimization problem. We demonstrate that the differentiable optimization layer we design has optimization feasibility during training and model representation capability. To the best of our knowledge, our work is the first visuomotor imitation learning framework that not only utilizes differentiable optimization but also demonstrates its efficacy in both simulation and real-world robotic manipulation tasks.
2) Safe and constrained action generation: In our approach, we formulate the optimization layer as a trajectory optimization problem. Through this approach, the policy generate constrained trajectories, enhancing the overall safety and smoothness of robot actions. Since the model incorporates position, velocity, and acceleration constraints during end-to-end training, it ensures robot safety without compromising performance when deploying the model as a real-time policy. Compared to the “black box” characteristics of neural networks, our approach can be described as a “gray box” that combines the safety and interpretability of optimization-based trajectory generation with the powerful representational capabilities of neural networks.
We evaluate LeTO both in simulations (Section IV) and on real robot tasks (Section V) quantitatively, and benchmark with diffusion policy [11], LSTM-GMM [12], and IBC [10]. In the simulated tests, LeTO achieves a success rate comparable to diffusion policy and better than LSTM-GMM and IBC, but the generated trajectories are of less uncertainty, higher quality, and smoother comparing with these methods. On the real robot, we implement constraints-critical manipulation tasks using LeTO, which highlights LeTO’s strengths.
@@ II-A Trajectory Optimization
Trajectory optimization is pivotal in generating safe and smooth trajectories for robots and remains an active area of research. CHOMP [13] approaches trajectory optimization by gradient techniques, focusing on optimizing the trade-off between safety and smoothness. TrajOpt [14] uses a sequential convex optimization to generate smooth and collision-free trajecotries and can use naive straight-line initializations that might be in collision. The minimum-snap trajectory optimization [15] uses piecewise polynomials to represent the trajectory and is optimized by quadratic programs. The work in [16] proposes a method for reformulating nondifferentiable collision avoidance constraints into smooth, differentiable constraints and enables real-time optimization-based trajectory planning.
In LeTO, we innovatively combine trajectory optimization with imitation learning by differentiable optimization, to generate smooth and constraint-compliant trajectories during policy rollout. Our model is trained end-to-end. During the training process, LeTO not only learns the policy from human demonstrations but also learns the appropriate trajectory optimization parameters suitable for the policy.
@@ II-B Imitation Learning
The basic way of imitation learning explicitly maps observations to actions [1, 17, 18, 19, 20, 21]. It can be trained using a regression loss. However, these policies are not ideal for capturing multi-modal distributions [10].
Prior works have aimed to represent multi-modal distributions by converting the regression into classification [22, 23, 24], using action discretization coupled with a multi-task action correction [25], using MDNs [12], and using implicit modeling that including energy-based model [10, 9] and diffusion model [2]. However, these methods overlook the critical aspects of safety for robot systems. For robotic systems, their behaviors need to meet specific constraints to ensure the safety of the system and the successful completion of tasks.
@@ II-C Differentiable Optimization in Robot Learning
Previous works utilize differentiable optimization for modeling discontinuous functions and dynamics [26, 27]. Additionally, some works focus on combining model-based methods with neural networks through differentiable optimization, such as using differentiable MPC [28] and koopman operator [29]. However, these methods are not suitable for tasks with high-dimensional observations, such as using camera images from multiple viewpoints.
@@
For robot navigation and obstacle avoidance, various end-to-end learning frameworks embed with differentiable optimization are proposed, such as using control barrier function [31], gradient-based correction [32], and stacking prediction, planning, and control in a differentiable way [33]. However, all these methods are not suited for performing manipulation tasks.
DiffTOP [34] is a method that integrate differentiable trajectory optimization as the policy representation to generate actions. In integrating imitation learning, DiffTOP focuses on capturing multi-modal distributions. However, it can not generating safe and constrained trajectories. In contrast, LeTO focuses on how the integration of differentiable trajectory optimization into the imitation learning framework enables the generation of trajectories that not only perform tasks effectively but also satisfy constraints.
@@
III Approach
In this paper, we assume access to an offline trajectories dataset of the task we want to do. The goal is to learn a policy from this dataset offline and can successfully do the task by running the policy online. The overview of LeTO is shown in Fig. 2. In this section, we will detail the design of the model.
@@ III-A Training Data Sampling
where $o$ represents the observations of the robot, such as images of the cameras with different views and robot states and $y\in\mathbb{R}^{D_{y}}$ represents the demonstrated action aligned with the observation $o$. $N_{d}$ is the total number of the demonstrated trajectories, $T_{d}$ is the total length of a trajectory, and $D_{y}$ is the dimension of action $y$.
Inspired by model predictive control and diffusion policy [2], in LeTO, we predict a action sequence at every time step. Moreover, we care about how to predict trajectories that satisfy constraints rather than action sequences simply fitting the human demonstration. Therefore, before training, we first sample many fixed-length sequences in following format from the dataset
@@
where $\mathbf{y}_{t}=\left[{y}_{t}^{\mathrm{T}},{y}_{t+1}^{\mathrm{T}},\ldots,{y}_{t+T_{p}-1}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{R}^{T_{p}D_{y}}$. $T_{s}$ is the length of each data sampling and $T_{p}$ is the length of the predicted action sequence, as shown in Fig. 3. The learning objective is minimizing the errors between the predicted $\hat{\mathbf{y}}^{*}_{t}$ and $\mathbf{y}_{t}$ (e.g. using MSE loss to assess the distance between them), as shown in Fig. 2. During policy rollout, the robot will execute the first $T_{a}$-step actions of $\hat{\mathbf{y}}^{*}_{t}$, and then repeat the same process based on new observations. This process is typically called receding horizon planning/control.
Continuous actions need to be constrained to ensure actions safety and smoothness while discrete actions do not. Therefore, we define a selection matrix $S$ to pick the continuous action vector $v$ that require to be constrained.
@@
As mentioned earlier, we focus on a velocity-based action space, so here, $v$ refers to the velocity at which the robot is being controlled and $D_{v}$ is the dimension of $v$.
@@ III-B Observation Encoder
The observation encoder maps the observations, like raw RGB images and robot states, into a latent embedding $e$ and is trained end-to-end with LeTO. To capture the sequential correlation of the task, we use a recurrent neural network (RNN) architecture, $E_{\theta}$, where the right subscript $\theta$ represent the parameters of the network. We choose long short-term memory (LSTM) as our RNN model. At each time step, $e_{t},h_{t+1}=E_{\theta}\left(o_{t},h_{t}\right),$ where $h$ is the hidden state. We define $e_{t}\in\mathbb{R}^{T_{p}D_{y}}$. More detailed reasons of setting the dimension as $T_{p}D_{y}$ can be seen at Section III-C.
@@ III-C Differentiable Trajectory Optimization Layer
The differentiable trajectory optimization (DTO) layer maps the embedding $e_{t}$ to a predicted action sequence $\hat{\mathbf{y}}_{t}=\left[\hat{y}_{t}^{\mathrm{T}},\hat{y}_{t+1}^{\mathrm{T}},\ldots,\hat{y}_{t+T_{p}-1}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{R}^{T_{p}D_{y}}$. In this section, we will show how to design this layer based on trajectory optimization, learn this layer during training, and integrate it with the whole neural network.
@@
$=\frac{1}{\Delta t}\left[\begin{matrix}-1&1&&&\\ &-1&1&&\\ &&\ddots&\ddots&\\ &&&-1&1\end{matrix}\right]\hat{\mathbf{v}}_{t}$ (7)
@@
$=\left[\begin{matrix}p_{t}\\ p_{t}\\ \vdots\\ p_{t}\end{matrix}\right]+\left[\begin{matrix}\Delta t&&&\\ \Delta t&\Delta t&\\ \vdots&\vdots&\ddots&\\ \Delta t&\Delta t&\ldots&\Delta t\end{matrix}\right]\hat{\mathbf{v}}_{t}$ (17)
@@
The sequence $\hat{\mathbf{p}}_{t},\hat{\mathbf{v}}_{t},\text{and}~{}\hat{\mathbf{a}}_{t}$ together constitute the trajectory generated by the policy. To optimize the trajectory, we must consider the relationship between $\hat{\mathbf{p}}_{t},\hat{\mathbf{v}}_{t},\text{and}~{}\hat{\mathbf{a}}_{t}$ and human demonstration, as well as the constraints of $\hat{\mathbf{p}}_{t},\hat{\mathbf{v}}_{t},\text{and}~{}\hat{\mathbf{a}}_{t}$. Based on this, the forward pass of the DTO layer is the following optimization problem:
@@
The coefficient $\alpha$ controls the smoothness of the trajectory. It can be observed in (19) that the larger the value of $\alpha$, the greater the cost associated with the sum of accelerations, leading the optimization problem to generate trajectories with a smaller sum of accelerations. $A_{\text{pos}}$ and $b_{\text{max,min}}$ define the convex constraints that the position must satisfy. $v_{\text{max,min}}$ and $a_{\text{max,min}}$ define the constraints for the velocity and acceleration of the trajectory. $\alpha,A_{\text{pos}},b_{\text{max,min}},v_{\text{max,min}}$ and $a_{\text{max,min}}$ are interpretable and model-based, so we do not need to learn these parameters during training. Instead, it is advisable to specify them according to the requirements of the task. On the contrary, $\mathbf{Q}\in\mathbb{R}^{T_{p}D_{y}}$ is a parameters that we need to learn during training. Next, we will demonstrate how to learn $\mathbf{Q}$ and its specific meaning in our trajectory optimization problem.
Optimization problem (20) is a quadratic program (QP). Therefore, the layer (20) can perform forward pass and backpropagation in batch form as long as it is always feasible in the training process [40]. The feasibility of (20) can be guaranteed by ensuring $\mathbf{Q}$ symmetric positive definite [41]. To achieve that, we utilize a Cholesky factorization
@@
If the constraint $\mathbf{b}_{\rm{min}}\leq\mathbf{A}_{\rm{pos}}\mathbf{p}_{t,0}\leq\mathbf{b}_{\rm{max}}$ is satisfied at each time step $t$, the optimization problem is always feasible, no matter how $\mathbf{L}$ and $e_{t}$ change. This means the optimization problem remains consistently feasible, which ensuring stable training process. $\mathbf{b}_{\rm{min}}\leq\mathbf{A}_{\rm{pos}}\mathbf{p}_{t,0}\leq\mathbf{b}_{\rm{max}}$ represents that the human demonstrations should satisfy the position constraints. Additionally, the human demonstrations do not need to satisfy the velocity and acceleration constraints strictly, because they have nothing to do with the feasibility of the optimization problem. Further more, for human demonstration, satisfying position constraints is much easier than satisfying the velocity and acceleration constraints.
Based on that $\mathbf{Q}$ is symmetric positive definite, the optimization problem (20) can be converted as the following least squares problem:
@@
DTO layer actually optimize a linear transformation under motion constraints. Since $\bar{\mathbf{L}}$ is a lower triangular matrix and it is fully ranked, the linear transformation here is
@@
Since the observation encoder is just a normal LSTM architecture, it is already has the power of representing the policy. Therefore, the whole model including the DTO layer has the power of fitting the human demonstration.

### v3

@@ Abstract
This paper introduces LeTO, a method for learning constrained visuomotor policy with differentiable trajectory optimization. Our approach integrates a differentiable optimization layer into the neural network. By formulating the optimization layer as a trajectory optimization problem, we enable the model to end-to-end generate actions in a safe and constraint-controlled fashion without extra modules. Our method allows for the introduction of constraint information during the training process, thereby balancing the training objectives of satisfying constraints, smoothing the trajectories, and minimizing errors with demonstrations. This “gray box” method marries optimization-based safety and interpretability with powerful representational abilities of neural networks. We quantitatively evaluate LeTO in simulation and in the real robot. The results demonstrate that LeTO performs well in both simulated and real-world tasks. In addition, it is capable of generating trajectories that are less uncertain, higher quality, and smoother compared to existing imitation learning methods. Therefore, it is shown that LeTO provides a practical example of how to achieve the integration of neural networks with trajectory optimization. We release our code at https://github.com/ZhengtongXu/LeTO.
Note to Practitioners
LeTO is driven by the goal of developing an imitation learning algorithm capable of generating safe and constraint-satisfying robotic behaviors. The idea of imitation learning is to enable the robot to learn from human demonstrations of certain tasks. Subsequently, the robot is able to autonomously perform the learned tasks on its own. Thanks to the powerful representational and fitting capabilities of neural networks, imitation learning can let robots perform complex manipulation tasks. However, neural networks often exhibit a certain level of uncertainty and lack theoretical safety guarantees. For robotic systems, it is crucial that robot behaviors meet specific constraints; otherwise, the system may not be sufficiently reliable. Therefore, we introduce LeTO, an approach that integrates trajectory optimization with neural networks to generate actions that not only achieve manipulation tasks, but also comply with constraints. This improves the interpretability, safety, and reliability of robot policies acquired through imitation learning, facilitating their deployment in scenarios with high safety requirements.
@@ I Introduction
Imitation learning [1] focuses on the derivation of robot policies from demonstrations. This process can be formulated as a supervised learning task, with the aim of learning the mapping between observations and robot actions. Through imitation learning, robots can perform highly complex and diverse tasks [2, 3, 4, 5, 6, 7, 8]. Moreover, recent studies have shown the feasibility of conducting imitation learning directly from manipulation videos [9, 10, 11].
However, ensuring the safety of robots is very important due to the interaction between robots and the real world. In contrast to optimization-based trajectory generation methods, imitation learning often exhibits greater uncertainty, which can lead to system instability, reduced robustness, and safety concerns. Hence, compared to other supervised learning problems, this poses novel and exceptional challenges in safety for imitation learning.
Existing research focus on different aspects and challenges of imitation learning, such as improving action accuracy by addressing compounding errors [12] and representing multi-modal distributions by implicit policy [13, 14, 15]. However, these methods overlook critical aspects of safety and smoothness of the generated trajectories. For robotic systems, their behaviors must meet specific constraints to ensure the safety of the system. Moreover, for certain tasks, robot actions also need to meet specific constraints to ensure their successful completion. In addition, adding explainable constraints for neural network-based algorithms, often called “black boxes”, is much harder than for traditional model-based methods.
@@
1) Differentiable optimization layer: We integrate a differentiable optimization layer into the neural network and formulate it as a trajectory optimization problem. We demonstrate that the proposed differentiable optimization layer is feasible for optimization during training and capable of effective model representation. To the best of our knowledge, our work is the first visuomotor imitation learning framework that not only utilizes differentiable optimization but also demonstrates its efficacy in both simulation and real-world robotic manipulation tasks.
2) Safe and constrained action generation: In our approach, we formulate the optimization layer as a trajectory optimization problem. Through this approach, the policy generates constrained trajectories, enhancing the overall safety and smoothness of robot actions. Since the model incorporates position, velocity, and acceleration constraints during end-to-end training, it ensures robot safety without compromising performance when deploying the model as a real-time policy. Compared to the “black box” characteristics of neural networks, our approach can be described as a “gray box” that combines the safety and interpretability of optimization-based trajectory generation with the powerful representational capabilities of neural networks. LeTO balances the objectives of skill learning and trajectory optimization with constraint guarantees.
In Section III, we will introduce the specific methodologies of LeTO, followed by the presentation of experimental results from simulations and real-world scenarios in Sections IV and V, respectively. Section VI will be the discussion section and will also propose future research questions that are promising and intriguing in our opinion.
@@ II-A Trajectory Optimization
Trajectory optimization is pivotal in generating safe and smooth trajectories for robots and remains an active area of research. CHOMP [16] approaches trajectory optimization by gradient techniques, focusing on optimizing the trade-off between safety and smoothness. TrajOpt [17] uses sequential convex optimization to generate smooth and collision-free trajecotries and can use naive straight-line initializations that might be in collision. The minimum-snap trajectory optimization [18] uses piecewise polynomials to represent the trajectory and is optimized by quadratic programs. The work in [19] proposes a method for reformulating nondifferentiable collision avoidance constraints into smooth, differentiable constraints and enables real-time optimization-based trajectory planning.
In LeTO, we innovatively combine trajectory optimization with imitation learning by differentiable optimization, to generate smooth and constraint-compliant trajectories during policy rollout. Our model is trained end-to-end. During the training process, LeTO not only learns the policy from human demonstrations but also learns the appropriate trajectory optimization parameters suitable for the policy. Moreover, we would like to clarify that our paper does not introduce a new trajectory generation algorithm. Rather, it focuses on presenting an imitation learning algorithm. The entire discussion within our paper is centered on this aspect of imitation learning.
@@ II-B Imitation Learning
The basic way of imitation learning explicitly maps observations to actions [1, 20, 21, 22, 23, 24]. It can be trained using regression loss. However, these policies are not ideal for capturing multi-modal distributions [14].
Previous works have aimed to represent multi-modal distributions by converting the regression into classification [25, 26, 27], using action discretization coupled with a multi-task action correction [28], using MDNs [29], and using implicit modeling that including energy-based model [14, 13] and diffusion model [15]. However, these methods overlook critical aspects of safety for robotic systems. For robotic systems, their behaviors must meet specific constraints to ensure the safety of the system and the successful completion of tasks.
@@ II-C Differentiable Optimization in Robot Learning
Previous works utilize differentiable optimization to model discontinuous functions and dynamics [30, 31]. Furthermore, some works focus on combining model-based methods with neural networks through differentiable optimization, such as using differentiable MPC [32] and koopman operator [33]. However, these methods are not suitable for tasks with high-dimensional observations, such as using camera images from multiple viewpoints.
@@
For robot navigation and obstacle avoidance, various end-to-end learning frameworks are proposed that are embedded with differentiable optimization, such as the use of the control barrier function [35], gradient-based correction [36], and stack prediction, planning, and control in a differentiable way [37]. However, all of these methods are not suitable for performing manipulation tasks.
DiffTOP [38] is a method that integrates differentiable trajectory optimization as the policy representation to generate actions. In integrating imitation learning, DiffTOP focuses on capturing multi-modal distributions. However, it cannot generate safe and constrained trajectories. In contrast, LeTO focuses on how the integration of differentiable trajectory optimization into the imitation learning framework enables the generation of trajectories that not only perform tasks effectively but also satisfy constraints.
III Approach
In this paper, we assume access to an offline trajectories dataset of the task we want to perform. The goal is to learn a policy from this dataset offline and can successfully carry out the task by running the policy online. The overview of LeTO is shown in Fig. 2. In this section, we will detail the design of the model.
@@ III-A Training Data Sampling
where $o$ represents the observations of the robot, such as images of the cameras with different views and robot states and $y\in\mathbb{R}^{D_{y}}$ represents the demonstrated action aligned with the observation $o$. $N_{d}$ is the total number of demonstrated trajectories, $T_{d}$ is the total length of a trajectory, and $D_{y}$ is the dimension of action $y$.
Inspired by model predictive control and diffusion policy [15], in LeTO, we predict an action sequence at each time step. Moreover, we care about how to predict trajectories that satisfy constraints rather than action sequences simply fitting the human demonstration. Therefore, before training, we first sample many fixed-length sequences in following format from the dataset
@@
where $\mathbf{y}_{t}=\left[{y}_{t}^{\mathrm{T}},{y}_{t+1}^{\mathrm{T}},\ldots,{y}_{t+T_{p}-1}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{R}^{T_{p}D_{y}}$. $T_{s}$ is the length of each data sampling and $T_{p}$ is the length of the predicted action sequence, as shown in Fig. 3. The learning objective is to minimize errors between the predicted $\hat{\mathbf{y}}^{*}_{t}$ and $\mathbf{y}_{t}$ (e.g. using MSE loss to assess the distance between them), as shown in Fig. 2. During policy rollout, the robot will execute the first $T_{a}$-step actions of $\hat{\mathbf{y}}^{*}_{t}$, and then repeat the same process based on new observations. This process is typically called receding horizon planning/control.
Continuous actions need to be constrained to ensure actions’ safety and smoothness, while discrete actions do not. Therefore, we define a selection matrix $S$ to pick the continuous action vector $v$ that must be constrained.
@@
As mentioned earlier, we focus on a velocity-based action space, so here $v$ refers to the velocity at which the robot is being controlled and $D_{v}$ is the dimension of $v$.
@@ III-B Observation Encoder
The observation encoder maps observations, such as raw RGB images and robot states, to a latent embedding $e$ and is trained end-to-end with LeTO. To capture the sequential correlation of the task, we use a recurrent neural network (RNN) architecture, $E_{\theta}$, where the right subscript $\theta$ represents the parameters of the network. We choose long short-term memory (LSTM) as our RNN model. At each time step, $e_{t},h_{t+1}=E_{\theta}\left(o_{t},h_{t}\right),$ where $h$ is the hidden state. We define $e_{t}\in\mathbb{R}^{T_{p}D_{y}}$. More detailed reasons of setting the dimension as $T_{p}D_{y}$ can be seen in Section III-C.
@@ III-C Differentiable Trajectory Optimization Layer
The differentiable trajectory optimization (DTO) layer maps the embedding $e_{t}$ to a predicted action sequence $\hat{\mathbf{y}}_{t}=\left[\hat{y}_{t}^{\mathrm{T}},\hat{y}_{t+1}^{\mathrm{T}},\ldots,\hat{y}_{t+T_{p}-1}^{\mathrm{T}}\right]^{\mathrm{T}}\in\mathbb{R}^{T_{p}D_{y}}$. In this section, we will show how to design this layer based on trajectory optimization, learn this layer during training, and integrate it with the entire neural network.
@@
$=\frac{1}{\Delta t}\left[\begin{matrix}-1&1&&&\\ &-1&1&&\\ &&\ddots&\ddots&\\ &&&-1&1\end{matrix}\right]\hat{\mathbf{v}}_{t}$
@@
$=\left[\begin{matrix}p_{t}\\ p_{t}\\ \vdots\\ p_{t}\end{matrix}\right]+\left[\begin{matrix}\Delta t&&&\\ \Delta t&\Delta t&\\ \vdots&\vdots&\ddots&\\ \Delta t&\Delta t&\ldots&\Delta t\end{matrix}\right]\hat{\mathbf{v}}_{t}$
@@
The sequence $\hat{\mathbf{p}}_{t},\hat{\mathbf{v}}_{t},\text{and}~{}\hat{\mathbf{a}}_{t}$ together constitutes the trajectory generated by the policy. To optimize the trajectory, we must consider the relationship between $\hat{\mathbf{p}}_{t},\hat{\mathbf{v}}_{t},\text{and}~{}\hat{\mathbf{a}}_{t}$ and human demonstration, as well as the constraints of $\hat{\mathbf{p}}_{t},\hat{\mathbf{v}}_{t},\text{and}~{}\hat{\mathbf{a}}_{t}$. Based on this, the forward pass of the DTO layer is the following optimization problem:
@@
The coefficient $\alpha$ controls the smoothness of the trajectory. It can be observed in (5) that the larger the value of $\alpha$, the higher the cost associated with the sum of accelerations, leading the optimization problem to generate trajectories with a smaller sum of accelerations. $A_{\text{pos}}$ and $b_{\text{max,min}}$ define the convex constraints that the position must satisfy. $v_{\text{max,min}}$ and $a_{\text{max,min}}$ define the constraints for the velocity and acceleration of the trajectory. $\alpha,A_{\text{pos}},b_{\text{max,min}},v_{\text{max,min}}$ and $a_{\text{max,min}}$ are interpretable and model-based, so we do not need to learn these parameters during training. Instead, it is advisable to specify them according to the requirements of the task. In contrast, $\mathbf{Q}\in\mathbb{R}^{T_{p}D_{y}}$ is a parameter that we need to learn during training. Next, we will demonstrate how to learn $\mathbf{Q}$ and its specific meaning in our trajectory optimization problem.
Optimization problem (6) is a quadratic program (QP). Therefore, the layer (6) can perform forward pass and backpropagation in batch form as long as it is always feasible in the training process [44]. The feasibility of (6) can be guaranteed by ensuring $\mathbf{Q}$ symmetric positive definite [45]. To achieve that, we use a Cholesky factorization
@@
If the constraint $\mathbf{b}_{\rm{min}}\leq\mathbf{A}_{\rm{pos}}\mathbf{p}_{t,0}\leq\mathbf{b}_{\rm{max}}$ is satisfied at each time step $t$, the optimization problem is always feasible, regardless of how $\mathbf{L}$ and $e_{t}$ change. This means that the optimization problem remains consistently feasible, which ensures a stable training process. $\mathbf{b}_{\rm{min}}\leq\mathbf{A}_{\rm{pos}}\mathbf{p}_{t,0}\leq\mathbf{b}_{\rm{max}}$ represents that the human demonstrations should satisfy the position constraints. Additionally, human demonstrations do not need to satisfy the velocity and acceleration constraints strictly, because they have nothing to do with the feasibility of the optimization problem. Further more, for human demonstration, satisfying position constraints is much easier than satisfying velocity and acceleration constraints.
Based on the fact that $\mathbf{Q}$ is symmetric positive definite, the optimization problem (6) can be converted to the following least squares problem:
@@
DTO layer actually optimizes a linear transformation under motion constraints. Since $\bar{\mathbf{L}}$ is a lower triangular matrix and it is fully ranked, the linear transformation here is
@@
Since the observation encoder is just a normal LSTM architecture, it is already has the power of representing the policy. Therefore, the whole model including the DTO layer has the power to fit the human demonstration.